# https://www.mitre.org/

### v1

@@ 4.4 Establishing Power Laws for Data Constrained Training
For all experiments we trained ScaleMAE ViT-Giant (ViT-G, 1.8B parameters) under identical hyperparameters, batch size, and optimization settings. This large backbone size was selected to ensure that learning was not limited by model capacity, only by dataset size. Training was conducted on a single H100 node (8 GPUs) for approximately 300k iterations per run. This controlled setup isolates the effect of dataset size while holding model capacity, compute and training dynamics constant. The resulting validation loss curves over training time are shown in figure 8. This design aims to test whether RS follows a stable data-scaling law within practical ranges and to quantify the exponent relevant for planning. Specifically, let $N$ denote the number of full-scene images and $L(N)$ denote the converged validation loss under a fixed architecture and schedule. We test the canonical power-law model $L(N)=A+BN^{-a}$, where $A$ is an asymptotic loss floor, $B$ is a scale factor, and $a$ is the data-scaling exponent. To obtain $a$, we perform log-log regression on $log(L)$ as a linear function of $log(N)$; we interpret slopes, variance, and residuals from this fit in section 5.3 and revisit their implications in section 6.
@@ 4.5 Establishing Power Laws for Parameter Constrained Training
To quantify how encoder capacity alone affects pretraining performance under fixed data and optimization dynamics, we constructed a parameter-constrained experiment that varies ViT backbone size, spanning increasing parameter counts across orders of magnitude (i.e., ViT-Base 86M [crossscale], ViT-Large 303M [scalemae], ViT-Huge 630M [scalingvits], ViT-Giant 1.8B [scalingvits]). To make the comparison with the data-constrained study explicit, we treat parameter scaling as a one-factor ablation: fix dataset composition, training schedule, and optimizer, and vary only backbone capacity. Let $P$ denote encoder parameters and let $L(P)$ denote the converged validation loss under a fixed training budget. Consistent with prior neural scaling formulations, we test whether $L(P)$ follows a power-law of the form $L(P)=A+BP^{-b}$, where $A$ approximates an asymptotic floor under the chosen data and supervision regime, $B$ is a scale factor, and $b$ indicates benefit from increased capacity. Once again, we perform linear regression on $log(L)$ vs. $log(P)$ to obtain a value for $b$. Refer to figure 10, section 5.4, and section 6 for loss curves, scaling analysis, and discussion of this experiment, respectively.
@@ 5 Results
Each subsection answers a guiding question: 5.1 (a: throughput vs. convergence), 5.2 (b: stable LR regimes), 5.3 (c: data scaling law) and 5.4 (d: parameter scaling law). We emphasize the practical takeaways and the conditions under which the observations hold.
@@ 5.1 Balancing Training Time and Compute Efficiency
Figure 5: Batch size scaling results show an increase in batch size does not result in increased compute efficiency during early-stage training, based on average loss over three trials. Based on the similar number of steps needed to reach convergence for the learning rates tested, we hypothesize that smaller batch sizes would maneuver this training configuration out of the saturated range and decrease forward pass compute efficiency.
Models trained with larger batches exhibited similar or occasionally slower, convergence compared to those with moderate batch sizes, suggesting that the critical batch size beyond which further increases yield diminishing returns in sample efficiency falls within or below this tested range. These findings are consistent with observations from language and vision foundation models, where critical batch size tends to depend more strongly on dataset scale than on model size [nlpscaling] [revistneuralscaling]. While larger batches may still offer benefits in late-stage training, when optimization stabilizes and gradient noise decreases, the current results suggest that early-stage training in geospatial foundation models remains batch-insensitive beyond moderate thresholds. This behavior reinforces that compute efficiency in large-scale pretraining should be sought primarily through data scaling and adaptive learning-rate scheduling, rather than indiscriminate batch size increases, which may improve throughput but introduce inefficiencies due to reduced gradient noise.
@@ 5.2 Optimal Learning Rate (LR)
Our fail fast preliminary warmup test demonstrated that higher learning rates (e.g., $1\times 10^{-3}$ and $3\times 10^{-3}$) exhibited significant instability, often diverging or oscillating during the stable phase of training, as shown in figure 6. Intermediate rates achieved stable convergence, confirming classic intuition that aggressive step sizes reduce optimization stability, and identifying a safe range allowing us to minimize the risk of catastrophic divergence without unnecessarily slowing training.
@@
Based on the results of our fail fast test, we graduated three learning rates to a full WSD cycle experiment; these results identify $1\times 10^{-4}$ as a functional learning rate for MTP with a ViT-L backbone while under WSD scheduling, as shown in figure 7. The finding aligns with prior scaling studies in language and vision where stable, moderate learning rates yield the best efficiency-performance balance across dataset sizes [revistneuralscaling].
Figure 7: Three learning rate candidates from our fail fast warmup experiment progress to a test of the full Warmup, Stable, Decay (WSD) schedule on Akupara-10k. Among the tested rates, $1\times 10^{-4}$ yields the most stable and efficient progress, maintaining monotonic loss decrease through the plateau and achieving the best final loss after decay, confirming that moderate base rates under WSD provide robust optimization, validate the predictive utility of the fail fast selection, and establish a reliable baseline for large-scale RS pretraining.
@@ 6 Discussion
Our findings demonstrate that large-scale RS foundation models follow predictable scaling dynamics, like those of related domains, and also exhibit domain-specific behaviors. The power-law established with an $\mathrm{R}^{2}$ fit of 0.941 between the dataset size and performance confirms that additional data consistently improves model quality, though the exponent $a=0.03$ indicates diminishing returns as datasets grow, consistent with trends observed in natural imagery and language domains [nlpscaling] [scalingvits]. Efficiency analyses reveal that compute gains from batch-size scaling plateau quickly and that training stability depends more on balanced learning rate schedules than raw throughput. These findings imply that further performance gains will come less from scaling compute and more from improving data diversity, label quality, and optimization strategy. Overall, RS models appear to operate in a data-limited, yet feature-bounded regime. Scaling continues to improve training progress, but only as it expands the domain of observed geospatial variation. Future scaling should therefore emphasize use of more feature-rich sensing modalities (e.g., SAR, multispectral, temporal), and geographically or contextually novel data, rather than simple quantity increases. The shallow parameter-scaling slope relative to data-scaling supports the view that RS pretraining is data-limited: capacity helps, but only insofar as it can absorb diverse, in-domain variation. In practice, prioritize expanding and domain coverage, then right-size to avoid undercapacity in rich regimes and overcapacity in narrow regimes.
@@ 6.1 Lessons learned and pitfalls
  • LR-induced instability: High base LRs under WSD produced oscillations and divergence; conservative LRs reduced instability.
  • Bounded spatial context: Increasing model size without expanding image context or modality yields marginal returns; invest in diversity and richer sensing (SAR, MSI, temporal).
@@ 6.2 Scope limitations
Our results reflect high-resolution Maxar imagery under ViT backbones and specific schedules; different sensors (e.g., SAR), data distributions, or architectural choices (e.g., CNN hybrids) may alter constants even if the qualitative trends persist.
@@ 7 Conclusion
This work establishes empirical scaling behavior for RS foundation models trained at peta-scale data volumes. We show that model loss decreases predictably with dataset size,while gains from larger batches or aggressive learning rates remain limited. These results suggest that sustainable progress in RS foundation modeling will hinge on use of more diverse data, not merely more compute. By quantifying how performance scales with data, this study provides a practical basis for planning future pretraining efforts and estimating the returns on additional data collection. Extending this framework to multi-modal, multi-temporal, and cross-sensor contexts offers a clear next step toward globally consistent geospatial representation learning. Future studies should compare training performance, as measured by optimization task loss, to downstream RS task performance to characterize the generalizability of our demonstrated scaling laws.

### v2

@@ 4.4 Establishing Power Laws for Data Constrained Training
For all experiments we trained ScaleMAE ViT-Giant (ViT-G, 1.8B parameters) under identical hyperparameters, batch size, and optimization settings. This large backbone size was selected to ensure that learning was not limited by model capacity, only by dataset size. Training was conducted on a single H100 node (8 GPUs) for approximately 300k iterations per run. This controlled setup isolates the effect of dataset size while holding model capacity, compute and training dynamics constant. The resulting validation loss curves over training time are shown in figure 8 in the results section.
This design aims to test whether RS follows a stable data-scaling law within practical ranges and to quantify the exponent relevant for planning. Specifically, let $N$ denote the number of full-scene images and $L(N)$ denote the converged validation loss under a fixed architecture and schedule. We test the canonical power-law model $L(N)=A+BN^{-a}$, where $A$ is an asymptotic loss floor, $B$ is a scale factor, and $a$ is the data-scaling exponent. To obtain $a$, we perform log-log regression on $log(L)$ as a linear function of $log(N)$; we interpret slopes, variance, and residuals from this fit in section 5.3 and revisit their implications in section 6.
@@ 4.5 Establishing Power Laws for Parameter Constrained Training
To quantify how encoder capacity alone affects pretraining performance under fixed data and optimization dynamics, we constructed a parameter-constrained experiment that varies ViT backbone size, spanning increasing parameter counts across orders of magnitude (i.e., ViT-Base 86M [10], ViT-Large 303M [7], ViT-Huge 630M [14], ViT-Giant 1.8B [14]). To make the comparison with the data-constrained study explicit, we treat parameter scaling as a one-factor ablation: fix dataset composition, training schedule, and optimizer, and vary only backbone capacity. Let $P$ denote encoder parameters and let $L(P)$ denote the converged validation loss under a fixed training budget. Consistent with prior neural scaling formulations, we test whether $L(P)$ follows a power-law of the form $L(P)=A+BP^{-b}$, where $A$ approximates an asymptotic floor under the chosen data and supervision regime, $B$ is a scale factor, and $b$ indicates benefit from increased capacity. Once again, we perform linear regression on $log(L)$ vs. $log(P)$ to obtain a value for $b$. Refer to figure 10, section 5.4, and section 6 for loss curves, scaling analysis, and discussion of this experiment, respectively.
@@ 5 Results
Each of the following subsections addresses one of our guiding questions, emphasize the practical takeaways, and the conditions under which the observations hold.
@@ 5.1 Balancing Training Time and Compute Efficiency
Figure 5: Batch size scaling results show an increase in batch size does not result in increased compute efficiency during early-stage training, based on average loss over three trials. Based on the similar number of steps needed to reach convergence for the LRs tested, we hypothesize that smaller batch sizes would maneuver this training configuration out of the saturated range and decrease forward pass compute efficiency.
Models trained with larger batches exhibited similar, or occasionally slower, convergence compared to those with moderate batch sizes, suggesting that the critical batch size beyond which further increases yield diminishing returns in sample efficiency falls within or below this tested range. These findings are consistent with observations from language and vision foundation models, where critical batch size tends to depend more strongly on dataset scale than on model size [6] [1]. While larger batches may still offer benefits in late-stage training, when optimization stabilizes and gradient noise decreases, the current results suggest that early-stage training in geospatial foundation models remains batch-insensitive beyond moderate thresholds. This behavior reinforces that compute efficiency in large-scale pretraining should be sought primarily through data scaling and adaptive learning-rate scheduling, rather than indiscriminate batch size increases, which may improve throughput but introduce inefficiencies due to reduced gradient noise.
@@ 5.2 Optimal Learning Rate (LR)
Our fail fast preliminary warmup test demonstrated that higher learning rates (e.g., $1\times 10^{-3}$ and $3\times 10^{-3}$) exhibited significant instability, often diverging or oscillating during the stable phase of training, as shown in figure 6. Intermediate rates achieved stable convergence, confirming classic intuition that aggressive step sizes reduce optimization stability, and identifying a safe range that allows us to minimize the risk of catastrophic divergence without unnecessarily slowing training.
@@
Based on the results of our fail fast test, we graduated three learning rates to a full WSD cycle experiment; these results identify $1\times 10^{-4}$ as a functional learning rate for MTP with a ViT-L backbone while under WSD scheduling, as shown in figure 7. The finding aligns with prior scaling studies in language and vision where stable, moderate learning rates yield the best efficiency-performance balance across dataset sizes [1].
Figure 7: Three learning rate candidates from our fail fast warmup experiment progress to a test of the full WSD schedule on Akupara-10k. Among the tested rates, $1\times 10^{-4}$ yields the most stable and efficient progress, maintaining monotonic loss decrease through the plateau and achieving the best final loss after decay, confirming that moderate base rates under WSD provide robust optimization, validate the predictive utility of the fail fast selection, and establish a reliable baseline for large-scale RS pretraining.
@@ 6 Discussion
Our findings demonstrate that large-scale RS foundation models follow predictable scaling dynamics, like those of related domains, and also exhibit domain-specific behaviors. The power-law established with an $\mathrm{R}^{2}$ fit of 0.941 between the dataset size and performance confirms that additional data consistently improves model quality, though the exponent $a=0.03$ indicates diminishing returns as datasets grow, consistent with trends observed in natural imagery and language domains [6] [14]. Efficiency analyses reveal that compute gains from batch-size scaling plateau quickly and that training stability depends more on balanced learning rate schedules than raw throughput. These findings imply that further performance gains will come less from scaling compute and more from improving data diversity, label quality, and optimization strategy. Overall, RS models appear to operate in a data-limited, yet feature-bounded regime. Scaling continues to improve training progress, but only as it expands the domain of observed geospatial variation. Future scaling should therefore emphasize use of more feature-rich sensing modalities (e.g., synthetic aperture radar (SAR), multispectral, temporal), and geographically or contextually novel data, rather than simple quantity increases. The shallow parameter-scaling slope in figure 11 relative to data-scaling in figure 9 supports the view that RS pretraining is data-limited: capacity helps, but only insofar as it can absorb diverse, in-domain variation. In practice, prioritize expanding and domain coverage, then right-size to avoid undercapacity in rich regimes and overcapacity in narrow regimes.
@@ 6.1 Lessons learned and pitfalls
  • LR-induced instability: High base LRs under WSD produced oscillations and divergence; conservative WSDs reduced instability.
  • Bounded spatial context: Increasing model size without expanding image context or modality yields marginal returns; invest in diversity and richer sensing (SAR, multispectral imagery (MSI), temporal).
@@ 6.2 Scope limitations
Our results reflect high-resolution imagery under ViT backbones and specific schedules; different sensors (e.g., SAR), data distributions, or architectural choices (e.g., CNN hybrids) may alter constants even if the qualitative trends persist.
@@ 7 Conclusion
This work establishes empirical scaling behavior for RS foundation models trained at peta-scale data volumes. We show that model loss decreases predictably with dataset size,while gains from larger batches or aggressive learning rates remain limited. These results suggest that sustainable progress in RS foundation modeling will hinge on use of more diverse data, not merely more compute. By quantifying how performance scales with data, this study provides a practical basis for planning future pretraining efforts and estimating the returns on additional data collection. Extending this framework to multimodal, multi-temporal, and cross-sensor contexts offers a clear next step toward globally consistent geospatial representation learning. Future studies should compare training performance, as measured by optimization task loss, to downstream RS task performance to characterize the generalizability of our demonstrated scaling laws.